a manufacture can sell x times. determine the total revenue function,profit function and level of output that will maximize profit
1
Expert's answer
2019-08-21T14:21:09-0400

A manufacturer can sell x items at the price of Rs. (330−x) each. The cost of producing x items is Rs. "x^2+10x-12"

The total revenue function


"R(x)=xp(x)=x(330-x)=330x-x^2"

The cost function


"C(x)=x^2+10x-12"

Profit


"P(x)=R(x)-C(x)"

"P(x)=330x-x^2- (x^2+10x-12)=-2x^2 +320x+12"

We need maximize the profit.

Find the first derivative with respect to x


"P'(x)=(-2x^2+320x+12)'=-4x+320"

Find the critical number(s)


"P'(x)=0=>-4x+320=0=>x=80"

First derivative test

If "x<80," then "P'(x)>0, P(x)" increases.

If "x>80," then "P'(x)<0, P(x)" decreases.


"P(80)=-2(80)^2+320(80)+12=12812"

The profit function "P(x)" has the local maximum with value of "12812" at "x=80."

Since "x=80" is the only critical number, then the profit function "P(x)" has the absolute maximum with value of "12812"

at "x=80."

The manufacturer can earn the maximum profit with value of "Rs. 12812" , if he sells "80" items.